What is the 20th term of the sequence: 1,1/3, . . . . . .,n +2/ 2n^2 + 3n - 2


1
Expert's answer
2021-08-19T15:30:38-0400

20th term,

"a_n=\frac{n+2}{2n^2+3n-2}\\
a_{20}=\frac{20+2}{2(20)^2+3(20)-2}\\
a_{20}=\frac{1}{39}\\"